# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v2

@@ 7 Discussion and Conclusion
In this paper, we focus on IRM, which is challenging to realize due to the difficulties of bi-level optimization. We conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to understand its approximation methods better.
The experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in calibration. Additionally, it was suggested that regularization penalties introduced for OOD generalization effectively achieve lower OOD ECE. However, on the other hand, it was observed that in more practical datasets (Diversity Shift), while the ECE of the model becomes lower, the OOD accuracy worsens. This indicates a trade-off between OOD ECE and accuracy, suggesting that carefully balancing these two metrics is important in aiming for OOD generalization.

### v3

@@ 7 Discussion and Conclusion
In this paper, we focus on IRM, which is challenging to implement due to the difficulties of bi-level optimization. We conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to understand its approximation methods better and measure how much the model learned environmental invariant features.
Our experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in calibration when we compared models with almost the same accuracy (Figure 2 and Table 2).
Another finding from our empirical results is that regularization penalties introduced for OOD generalization effectively achieve lower ECE. However, in practical datasets (under diversity shift), we observe a trade-off between OOD ECE and accuracy. Although IB-IRM mitigates this tradeoff compared to IRMv1 and other variants; carefully balancing these two metrics is important in aiming for OOD generalization.